OpenStudy (anonymous):

n the physics of Optics (the study of light) the magnitude (brightness) M of a certain polarized light is given by the equation M=k sin 2θ,…

OpenStudy (anonymous):

sin 42 = x/19

OpenStudy (mertsj):

Ok. aajugdar wants to help you now.

OpenStudy (mertsj):

There you go. @aajugdar

OpenStudy (anonymous):

I don't get it @aajugdar ?

OpenStudy (anonymous):

okay letme draw

OpenStudy (anonymous):



OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

isn't it right?

OpenStudy (anonymous):

I don't see that as a choice?

OpenStudy (anonymous):

answer is 12.7 ~

OpenStudy (anonymous):

oh come on sin 42 = x/19 x= 19 * sin42 19*0.6691 = 12.7

OpenStudy (anonymous):

got it now?

OpenStudy (anonymous):

ohhhhhh k npw it makes sense

OpenStudy (anonymous):

ofcrs opposite side would be tree because u wanna calculate height from ground where cat is stuck